How high can the river rise in 5 days if the first day the river rose 3 inches & each day after it rose twice as much as the previous day ?

Answer by ikleyn(52818)

.
How high can the river rise in 5 days if the first day the river rose 3 inches & each day after it rose twice as much as the previous day ?
-----------------------------------------------------------

The sum of the five first terms of a Geom. progression with the 1-st term 3 and common difference 2:

3 + 2*3 + 4*3 + 8*3 + 16*3 = 3*(1 + 2 + 4 + 8 + 16) = 3*31 = 93 in.